5-4x 8
---- = -
2-9x 7

Answer by checkley77(12569)

5-4x 8
---- = - CROSS MULTIPLY.
2-9x 7
8(2-9X)=7(5-4X)
16-72X=35-28X
-72X+28X=35-16
-44X=19
X=19/-44
X=-19/44 ANSWER.